OpenStudy (anonymous):

Nick's restaurant bill totaled sixty-four dollars and twenty cents. Show how to estimate fifteen percent of the tip. Show and explain each step

OpenStudy (anonymous):

64.2 * 15/100 = approx 64*15/100 = 960/100=9.6